# Corresponding author(s)

### v1

@@ 2.1 Instruction-tuning Data Collection
Currently, the resources for English instruction-tuning datasets are relatively abundant, while datasets for non-English languages generally face challenges in terms of small scale or quality issues. To construct instruction-tuning datasets for specific non-English languages, existing methods primarily include: 1) manual annotation[6], which is precise but time-consuming and labor-intensive, resulting in high costs in practical applications; 2) translating English open-source datasets into specific languages, yet this approach may accumulate errors due to translation inaccuracies; 3) utilizing LLMs for automatic generation [7, 8, 13, 14, 15], but this method is limited by the model’s internal capabilities and may lead to hallucinations.

### v2

@@ 2.1 Instruction-tuning Data Collection
Currently, the resources for English instruction-tuning datasets are relatively abundant, while datasets for non-English languages generally face challenges in terms of small scale or quality issues. To construct instruction-tuning datasets for specific non-English languages, existing methods primarily include: 1) manual annotation[6], which is precise but time-consuming and labor-intensive, resulting in high costs in practical applications; 2) translating English open-source datasets into specific languages[10], yet this approach may accumulate errors due to translation inaccuracies; 3) utilizing LLMs for automatic generation [7, 8, 13, 14, 15], but this method is limited by the model’s internal capabilities and may lead to hallucinations.